A family of four members. The father is 2 years older than the mother. The sister is 2 years older than the brother. The sum of all the present ages is 64. Three years ago. The sum of their agrs was 53. How old is each of them now?
Answer by stanbon(75887)

A family of four members. The father is 2 years older than the mother. The sister is 2 years older than the brother. The sum of all the present ages is 64. Three years ago. The sum of their agrs was 53. How old is each of them now?
------------
Equations:
f = m + 2
s = b + 2
f + m + b + s = 64
f + m + b + s -12 = 53
--------
Modify to get:
f = m + 2
s = b + 2
f + m + b + s = 64
f + m + b + s = 65
----
Note: The last two equation are contradictory.
----
No solution.
Cheers,
Stan H.